OpenStudy (anonymous):

find the volume of the box .volume is 160 find lenght x+2 width x height 2 so am guesiing that you will write it 160=(x+2)*x*2

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

You guess correctly.

OpenStudy (anonymous):

so know i am a lil confused on the next step

OpenStudy (anonymous):

this is so far what i got and the part that am stuck on x^2+2x*2=160

OpenStudy (anonymous):

(x^2+2x)2=160 (x^2+2x)=80 Factor out and x and solve it twice.

OpenStudy (anonymous):

am not sure what you mean?

OpenStudy (anonymous):

Do you understand how I got from the top line to the bottom one?

OpenStudy (anonymous):

yes you divide by two

OpenStudy (anonymous):

Now take the x^2+2x and factor out an x: x(...)=80

OpenStudy (anonymous):

ok so I get x(x+2)=80 wat do i do after

OpenStudy (anonymous):

Whoops. I'm sorry. Take that last step back. x^2+2x=80 => x^2+2x-80=0 Factor that equation.

OpenStudy (anonymous):

so know i got(x+10)(x-8)

OpenStudy (anonymous):

Well no. That is an expression. You have an equation: (x+10)(x-8)=0 Solve each of those for 0 and you get two answers.